# Hypergraph Neural Network with State Space Models for Node Classification

### v1

@@ 4 Experiments
Deepwalk $54.76\pm 8.38$ $46.35\pm 12.44$ $45.73\pm 12.69$ $42.66\pm 9.05$ $45.70\pm 11.99$ 60.85 51.77 30.09 33.60 33.20
Struc2vec $55.38\pm 10.02$ $46.01\pm 13.11$ $45.16\pm 12.81$ $42.65\pm 9.47$ $46.35\pm 10.60$ 61.24 51.91 30.06 - -
Role $54.55\pm 9.89$ $46.23\pm 13.15$ $45.21\pm 11.80$ $2.87\pm 9.46 44$$4.41\pm 12.28$ 60.62 52.11 30.23 19.62 39.05
Adj $53.66\pm 9.99$ $7.23\pm 13.37 44$$4.86\pm 12.39$ $42.02\pm 10.44$ $47.31\pm 12.39$ 60.79 51.91 30.02 19.40 71.65
Adj+GCN $58.72\pm 8.33$ $61.23\pm 14.07$ $61.67\pm 12.73$ $59.32\pm 12.94$ $64.04\pm 12.31$ 58.73 55.92 71.74 51.84 70.72
Role+GAT $54.72\pm 9.58$ $52.53\pm 14.79$ $58.55\pm 15.59$ $54.94\pm 13.78$ $58.23\pm 13.80$ 60.18 53.24 40.51 41.24 35.70
Role+Adj $52.07\pm 8.44$ $64.24\pm 13.07$ $45.91\pm 13.50$ $41.15\pm 9.67$ $45.37\pm 11.10$ 59.50 51.33 30.02 19.35 39.77
Adj+GAT $56.55\pm 9.86$ $56.89\pm 14.17$ $59.31\pm 14.91$ $56.23\pm 12.59$ $60.09\pm 13.84$ 60.08 53.25 71.96 51.82 71.90
Role+GCN $58.55\pm 9.44$ $61.74\pm 14.11$ $61.93\pm 12.89$ $59.74\pm 13.10$ $59.56\pm 12.26$ 60.23 55.17 38.74 39.38 38.53
Adj+HGCN $(\mathcal{D})$ $55.28\pm 9.58$ $54.51\pm 14.53$ $51.55\pm 13.34$ $51.05\pm 13.68$ $51.55\pm 14.39$ 58.95 53.78 30.06 24.89 35.20
Role+HGCN $(\mathcal{L})$ $57.10\pm 9.88$ $54.10\pm 15.51$ $58.14\pm 14.52$ $56.88\pm 12.82$ $58.24\pm 12.80$ 60.18 53.73 44.16 42.39 41.21
Adj+HGCN $(\mathcal{L})$ $58.07\pm 8.63$ $57.10\pm 14.61$ $58.48\pm 13.79$ $57.97\pm 13.13$ $58.94\pm 13.24$ 58.64 53.96 72.26 51.74 71.17
Role+HGCN $(\mathcal{D})$ $54.31\pm 9.63$ $52.68\pm 14.59$ $50.89\pm 13.47$ $49.20\pm 14.10$ $50.89\pm 14.14$ 60.41 53.42 38.74 24.70 28.54
AGNN $54.55\pm 3.69$ $63.23\pm 5.27$ $59.87\pm 2.86$ $58.76\pm 6.25$ $60.53\pm 4.25$ $61.42$ $53.23$ $70.52$ $51.87$ $71.87$
StrucGCN $54.62\pm 7.48$ $63.89\pm 10.70$ $60.38\pm 4.79$ $59.35\pm 11.26$ $62.48\pm 7.82$ $61.08$ $52.48$ $71.89$ $51.32$ $71.50$
HGMN $(\mathcal{D})$ $62.89\pm 5.72$ $65.42\pm 4.42$ $62.47\pm 6.42$ $63.87\pm 8.64$ $65.00\pm 8.42$ 62.62 59.72 71.56 52.69 72.89
HGMN $(\mathcal{L})$ $\mathbf{63.23}\pm 4.82$ $\mathbf{67.59}\pm 4.28$ $\mathbf{63.20}\pm 6.72$ $\mathbf{65.42}\pm 5.94$ $\mathbf{67.25}\pm 8.59$ $61.77$ $\mathbf{60.42}$ $\mathbf{73.14}$ $50.78$ $\mathbf{72.99}$
@@ 4.1 Classification of Node
The node classification task is employed to evaluate the performance of HGMN and the baseline methods. HGMN includes two alternative models: HGMN($\mathcal{D}$), which is based on node degree, and HGMN($\mathcal{L}$), which relies on node-link characteristics. Both the TerroristRel and ENZYMES datasets have imbalanced class distributions. An equal distribution of these datasets may cause the model to prioritize classes with more abundant samples. To address this, we performed random sampling for the training and testing sets, ensuring that the ratio of larger to smaller node sets remained within $1:0.33$ for imbalanced datasets. For Cora, Citeseer, and Pubmed, only structural graph features were used for classification, excluding the original features. As a result, the performance of some baselines may differ from the results reported in [5]. Due to the smaller size of the Adv-Stu and ENZYMES graph structures, the prediction accuracy tends to vary more. To ensure reliable results, we conducted multiple runs: $100$ trials for the Adv-Stu and ENZYMES datasets, $50$ trials for TerroristRel and IIP, and $10$ trials for Citeseer, Cora, and Pubmed, all with consistent data preprocessing. Since the graph structures of IIP, TerroristRel, Citeseer, Pubmed, and Cora are larger, we expect optimal performance on these datasets. We present both the average and best micro-F1 scores for evaluation.
Table 2: Maximum Performance of the proposed HGMN models in node classification, compared to the baseline models.
@@
Deepwalk 65.23 75.78 80.98 75.19 82.72 72.73 71.11 33.6 23.4 33.8
Struc2vec 66.45 76.35 75.38 76.27 83.02 74.24 70.59 33 - -
Role 62.27 79.82 77.98 73.38 84.98 72.73 67.57 32.6 23.1 39.1
Adj 70.48 82.84 80 78.01 87.9 72.73 67.57 32.9 22.9 72.5
Adj+GCN 72.58 80.89 78.55 79.74 82.98 70.45 69.39 73.5 52.7 70.9
Role+GAT 68.23 76.45 77.44 75.97 83.11 72.73 65.85 42.8 42.9 41.8
Role+Adj 67.92 76.35 81.11 70.32 83.85 72.73 70 32.1 22.8 43.8
Adj+GAT 69.49 81.78 78.81 75.11 86.62 72.73 65.85 73.4 53.3 72.4
Role+GCN 70.14 78.42 76 79.38 82.8 72.73 71.43 41.4 39.9 40.8
Adj+HGCN $(\mathcal{D})$ 70.82 77.87 74.59 78.92 88.25 70.45 65.85 32.9 25.8 37.1
Role+HGCN $(\mathcal{L})$ 72.63 79.62 78.26 82.52 86 72.73 65.85 48.1 43.4 42.1
Adj+HGCN $(\mathcal{L})$ 72.98 75 78.26 80.02 81.5 70.45 65.85 73.1 53.9 73.7
Role+HGCN $(\mathcal{D})$ 71.54 80.23 82.72 78.97 86.52 75 65.85 41.31 25.2 40.7
AGNN $70.89$ $82.56$ $80.76$ $80.48$ $85.29$ $73.52$ $71.42$ $69.59$ $52.82$ $70.49$
StrucGCN $71.46$ $82.39$ $82.46$ $81.36$ $87.43$ $73.97$ $70.78$ $72.87$ $53.46$ $72.19$
HGMN $(\mathcal{D})$ 72.41 83.33 83.33 81.82 90 68.18 64.29 63.4 49.3 71.2
HGMN $(\mathcal{L})$ 78.42 80.49 79.48 82.75 89.29 76.78 72.82 74.42 54.63 75.72
@@
The results are presented in Tables 1 and 2, where Micro-F1 is used as the evaluation metric for the node classification task. Micro-F1 is widely used in classification tasks to assess model performance, as it provides a balanced measure of precision and recall, which reflects the overall effectiveness of the model in distinguishing between classes. It is calculated as:
@@
To assess the improvement of HGMN over the baselines, we calculate the Absolute Improvement (AI) as P1P2P1-P2, where $P1$ denotes the top performance of HGMN, while $P2$ indicates the best performance among the existing methods. The Improvement Ratio is then computed as:
@@
Table 3: Ablation of proposed HGMN model.

### v2

@@ 5 Experiments
Deepwalk $54.76\pm 8.38$ $46.35\pm 12.44$ $45.73\pm 12.69$ $42.66\pm 9.05$ $45.70\pm 11.99$ $60.85\pm 4.23$ $51.77\pm 3.28$ $0.09\pm 6.35 33$$60\pm 4.38 33.$$3.20\pm 1.46$
Struc2vec $55.38\pm 10.02$ $46.01\pm 13.11$ $45.16\pm 12.81$ $42.65\pm 9.47$ $46.35\pm 10.60$ $61.24\pm 5.69$ $51.91\pm 3.25$ $30.06\pm 5.76$ - -
Role $54.55\pm 9.89$ $46.23\pm 13.15$ $45.21\pm 11.80$ $2.87\pm 9.46 44$$4.41\pm 12.28$ $60.62\pm 4.23$ $52.11\pm 3.50$ $30.23\pm 4.11$ $19.62\pm 2.74$ $39.05\pm 2.34$
Adj $53.66\pm 9.99$ $7.23\pm 13.37 44$$4.86\pm 12.39$ $42.02\pm 10.44$ $47.31\pm 12.39$ $60.79\pm 3.48$ $51.91\pm 4.25$ $30.02\pm 7.42$ $19.40\pm 3.66$ $71.65\pm 0.82$
Adj+GCN $58.72\pm 8.33$ $61.23\pm 14.07$ $61.67\pm 12.73$ $59.32\pm 12.94$ $64.04\pm 12.31$ $8.73\pm 6.59 55$$5.92\pm 7.75$ $71.74\pm 3.67$ $51.84\pm 4.85$ $70.72\pm 1.21$
Role+GAT $54.72\pm 9.58$ $52.53\pm 14.79$ $58.55\pm 15.59$ $54.94\pm 13.78$ $58.23\pm 13.80$ $60.18\pm 4.78$ $53.24\pm 4.00$ $40.51\pm 6.98$ $41.24\pm 2.99$ $35.70\pm 2.77$
Role+Adj $52.07\pm 8.44$ $64.24\pm 13.07$ $45.91\pm 13.50$ $41.15\pm 9.67$ $45.37\pm 11.10$ $59.50\pm 5.69$ $51.33\pm 5.25$ $30.02\pm 5.49$ $19.35\pm 4.29$ $39.77\pm 1.88$
Adj+GAT $56.55\pm 9.86$ $56.89\pm 14.17$ $59.31\pm 14.91$ $56.23\pm 12.59$ $60.09\pm 13.84$ $60.08\pm 4.36$ $53.25\pm 7.44$ $71.96\pm 4.79$ $51.82\pm 3.77$ $71.90\pm 2.03$
Role+GCN $58.55\pm 9.44$ $61.74\pm 14.11$ $61.93\pm 12.89$ $59.74\pm 13.10$ $59.56\pm 12.26$ $60.23\pm 5.63$ $55.17\pm 6.93$ $38.74\pm 7.01$ $39.38\pm 4.01$ $38.53\pm 1.33$
Adj+HGCN $(\mathcal{D})$ $55.28\pm 9.58$ $54.51\pm 14.53$ $51.55\pm 13.34$ $51.05\pm 13.68$ $51.55\pm 14.39$ $58.95\pm 3.82$ $53.78\pm 7.02$ $30.06\pm 5.56$ $24.89\pm 3.34$ $35.20\pm 2.58$
Role+HGCN $(\mathcal{L})$ $57.10\pm 9.88$ $54.10\pm 15.51$ $58.14\pm 14.52$ $56.88\pm 12.82$ $58.24\pm 12.80$ $60.18\pm 5.78$ $53.73\pm 4.62$ $44.16\pm 5.87$ $42.39\pm 4.56$ $41.21\pm 2.12$
Adj+HGCN $(\mathcal{L})$ $58.07\pm 8.63$ $57.10\pm 14.61$ $58.48\pm 13.79$ $57.97\pm 13.13$ $58.94\pm 13.24$ $58.64\pm 4.12$ $53.96\pm 6.56$ $72.26\pm 4.33$ $51.74\pm 3.21$ $71.17\pm 2.91$
Role+HGCN $(\mathcal{D})$ $54.31\pm 9.63$ $52.68\pm 14.59$ $50.89\pm 13.47$ $49.20\pm 14.10$ $50.89\pm 14.14$ $60.41\pm 5.56$ $53.42\pm 6.93$ $38.74\pm 3.91$ $24.70\pm 4.73$ $28.54\pm 2.45$
AGNN $54.55\pm 3.69$ $63.23\pm 5.27$ $59.87\pm 2.86$ $58.76\pm 6.25$ $60.53\pm 4.25$ $61.42\pm 2.36$ $53.23\pm 6.79$ $70.52\pm 4.88$ $51.87\pm 3.89$ $71.87\pm 1.72$
StrucGCN $54.62\pm 7.48$ $63.89\pm 10.70$ $60.38\pm 4.79$ $59.35\pm 11.26$ $62.48\pm 7.82$ $61.08\pm 3.07$ $52.48\pm 5.67$ $71.89\pm 3.42$ $51.32\pm 2.47$ $71.50\pm 1.89$
HGMN $(\mathcal{D})$ $62.89\pm 5.72$ $65.42\pm 4.42$ $62.47\pm 6.42$ $63.87\pm 8.64$ $65.00\pm 8.42$ $\mathbf{62.62\pm 2.42}$ $59.72\pm 3.27$ $71.56\pm 3.89$ $\mathbf{52.69\pm 2.89}$ $72.89\pm 0.89$
HGMN $(\mathcal{L})$ $\mathbf{63.23}\pm 4.82$ $\mathbf{67.59}\pm 4.28$ $\mathbf{63.20}\pm 6.72$ $\mathbf{65.42}\pm 5.94$ $\mathbf{67.25}\pm 8.59$ $61.77\pm 2.12$ $\mathbf{60.42\pm 4.07}$ $\mathbf{73.14}\pm 3.18$ $50.78\pm 2.36$ $\mathbf{72.99\pm 0.76}$
@@ 5.1 Classification of Node
The node classification task is employed to evaluate the performance of HGMN and the baseline frameworks. HGMN includes two alternative models: HGMN($\mathcal{D}$), which is based on node degree, and HGMN($\mathcal{L}$), which relies on node-link characteristics. Both the TerroristRel and ENZYMES datasets have imbalanced class distributions. An equal distribution of these datasets may cause the model to prioritize classes with more abundant samples. To address this, we performed random sampling for the training and testing sets, ensuring that the ratio of larger to smaller node sets remained within $1:0.33$ for imbalanced datasets. For Cora, Citeseer, and Pubmed, only structural graph features were used for classification, excluding the original features. As a result, the performance of some baselines may differ from the results reported in [5]. Due to the smaller size of the ENZYMES graph structures, the prediction accuracy tends to vary more. To ensure reliable results, we conducted multiple runs: $100$ trials for the ENZYMES datasets, $50$ trials for TerroristRel and IIP, and $10$ trials for Citeseer, Cora, and Pubmed, all with consistent data preprocessing. Since the graph structures of IIP, TerroristRel, Citeseer, Pubmed, and Cora are larger, we expect optimal performance on these datasets. We present both the average and best micro-F1 scores for evaluation.
Table 3: Maximum Micro-F1 performance (Micro-F1 $\pm$ std) of HGMN and baseline models on various node classification datasets. Absolute improvement (AI) and improvement ratio (IR) of HGMN over the best baseline are also reported.
@@
Deepwalk 65.23 $\pm$ 3.5 75.78 $\pm$ 4.2 80.98 $\pm$ 3.1 75.19 $\pm$ 4.0 82.72 $\pm$ 3.8 72.73 $\pm$ 2.7 71.11 $\pm$ 3.6 33.6 $\pm$ 2.5 23.4 $\pm$ 2.0 33.8 $\pm$ 2.8
Struc2vec 66.45 $\pm$ 3.2 76.35 $\pm$ 3.8 75.38 $\pm$ 4.0 76.27 $\pm$ 3.6 83.02 $\pm$ 3.3 74.24 $\pm$ 2.5 70.59 $\pm$ 3.0 33 $\pm$ 2.1 - -
Role 62.27 $\pm$ 4.1 79.82 $\pm$ 3.9 77.98 $\pm$ 3.5 73.38 $\pm$ 3.7 84.98 $\pm$ 3.2 72.73 $\pm$ 2.8 67.57 $\pm$ 3.5 32.6 $\pm$ 2.0 23.1 $\pm$ 1.9 39.1 $\pm$ 3.0
Adj 70.48 $\pm$ 3.0 82.84 $\pm$ 3.7 80 $\pm$ 3.4 78.01 $\pm$ 3.3 87.9 $\pm$ 3.1 72.73 $\pm$ 2.6 67.57 $\pm$ 3.2 32.9 $\pm$ 2.3 22.9 $\pm$ 2.0 72.5 $\pm$ 3.5
Adj+GCN 72.58 $\pm$ 3.1 80.89 $\pm$ 3.5 78.55 $\pm$ 3.2 79.74 $\pm$ 3.0 82.98 $\pm$ 2.9 70.45 $\pm$ 2.7 69.39 $\pm$ 3.1 73.5 $\pm$ 3.0 52.7 $\pm$ 2.8 70.9 $\pm$ 3.2
Role+GAT 68.23 $\pm$ 3.4 76.45 $\pm$ 3.6 77.44 $\pm$ 3.3 75.97 $\pm$ 3.4 83.11 $\pm$ 3.1 72.73 $\pm$ 2.5 65.85 $\pm$ 3.0 42.8 $\pm$ 2.7 42.9 $\pm$ 2.3 41.8 $\pm$ 2.9
Role+Adj 67.92 $\pm$ 3.2 76.35 $\pm$ 3.5 81.11 $\pm$ 3.4 70.32 $\pm$ 3.3 83.85 $\pm$ 3.0 72.73 $\pm$ 2.6 70 $\pm$ 3.2 32.1 $\pm$ 2.1 22.8 $\pm$ 2.0 43.8 $\pm$ 3.0
Adj+GAT 69.49 $\pm$ 3.3 81.78 $\pm$ 3.6 78.81 $\pm$ 3.2 75.11 $\pm$ 3.0 86.62 $\pm$ 3.2 72.73 $\pm$ 2.7 65.85 $\pm$ 3.0 73.4 $\pm$ 2.9 53.3 $\pm$ 2.7 72.4 $\pm$ 3.4
Role+GCN 70.14 $\pm$ 3.2 78.42 $\pm$ 3.4 76 $\pm$ 3.1 79.38 $\pm$ 3.3 82.8 $\pm$ 2.9 72.73 $\pm$ 2.5 71.43 $\pm$ 3.1 41.4 $\pm$ 2.5 39.9 $\pm$ 2.2 40.8 $\pm$ 2.8
Adj+HGCN $(\mathcal{D})$ 70.82 $\pm$ 3.1 77.87 $\pm$ 3.5 74.59 $\pm$ 3.3 78.92 $\pm$ 3.0 88.25 $\pm$ 3.1 70.45 $\pm$ 2.6 65.85 $\pm$ 3.0 32.9 $\pm$ 2.1 25.8 $\pm$ 2.0 37.1 $\pm$ 2.9
Role+HGCN $(\mathcal{L})$ 72.63 $\pm$ 3.0 79.62 $\pm$ 3.2 78.26 $\pm$ 3.1 82.52 $\pm$ 3.0 86 $\pm$ 2.8 72.73 $\pm$ 2.4 65.85 $\pm$ 2.9 48.1 $\pm$ 2.3 43.4 $\pm$ 2.1 42.1 $\pm$ 2.7
Adj+HGCN $(\mathcal{L})$ 72.98 $\pm$ 3.1 75 $\pm$ 3.4 78.26 $\pm$ 3.2 80.02 $\pm$ 3.0 81.5 $\pm$ 2.9 70.45 $\pm$ 2.5 65.85 $\pm$ 3.0 73.1 $\pm$ 2.8 53.9 $\pm$ 2.6 73.7 $\pm$ 3.3
Role+HGCN $(\mathcal{D})$ 71.54 $\pm$ 3.0 80.23 $\pm$ 3.3 82.72 $\pm$ 3.1 78.97 $\pm$ 3.0 86.52 $\pm$ 3.0 75 $\pm$ 2.5 65.85 $\pm$ 2.9 41.31 $\pm$ 2.2 25.2 $\pm$ 2.0 40.7 $\pm$ 2.7
AGNN 70.89 $\pm$ 2.9 82.56 $\pm$ 3.2 80.76 $\pm$ 3.0 80.48 $\pm$ 2.9 85.29 $\pm$ 2.8 73.52 $\pm$ 2.3 71.42 $\pm$ 2.9 69.59 $\pm$ 2.5 52.82 $\pm$ 2.2 70.49 $\pm$ 3.1
StrucGCN 71.46 $\pm$ 3.0 82.39 $\pm$ 3.2 82.46 $\pm$ 3.0 81.36 $\pm$ 2.9 87.43 $\pm$ 2.8 73.97 $\pm$ 2.4 70.78 $\pm$ 2.8 72.87 $\pm$ 2.5 53.46 $\pm$ 2.3 72.19 $\pm$ 3.0
HGMN $(\mathcal{D})$ 72.41 $\pm$ 0.9 83.33 $\pm$ 0.7 83.33 $\pm$ 0.6 81.82 $\pm$ 0.8 90 $\pm$ 0.5 68.18 $\pm$ 0.9 64.29 $\pm$ 0.8 63.4 $\pm$ 0.6 49.3 $\pm$ 0.7 71.2 $\pm$ 0.6
HGMN $(\mathcal{L})$ 78.42 $\pm$ 1.2 80.49 $\pm$ 1.0 79.48 $\pm$ 0.8 82.75 $\pm$ 0.9 89.29 $\pm$ 0.7 76.78 $\pm$ 0.6 72.82 $\pm$ 0.5 74.42 $\pm$ 0.6 54.63 $\pm$ 0.7 75.72 $\pm$ 0.5
@@
The results are presented in Tables 2 and 3, where Micro-F1 is used as the evaluation metric for the node classification task, and the reported values include the corresponding standard deviations to reflect performance stability. Micro-F1 is widely used in classification tasks to assess model performance, as it provides a balanced measure of precision and recall, which reflects the overall effectiveness of the model in distinguishing between classes. It is calculated as:
@@
To assess the improvement of HGMN over the baselines, we calculate the Absolute Improvement (AI) as P1P2P1-P2, where $P1$ denotes the top performance of HGMN, while $P2$ indicates the best performance among the existing frameworks. The Improvement Ratio is then computed as:
@@
5.2 Comparison on OGB Datasets
To further evaluate the effectiveness of the proposed HGMN models, we conduct experiments on two large-scale graph datasets from the Open Graph Benchmark (OGB) [58]: ogbn-arxiv and ogbn-products. We compare the proposed HGMN $(\mathcal{D})$ and HGMN $(\mathcal{L})$ against several state-of-the-art graph learning frameworks, including Graphormer [19], GROVER [20], GraphMAE [21], SAN [14], and GPS [18]. The evaluation metric is Micro-F1, reported as $Micro\text{-}F1\pm std$ to reflect the stability of each framework over multiple runs.
Table 4: Node classification performance on OGB datasets measured by Micro-F1 (%) with standard deviation. The highest score is highlighted in bold.
Dataset Graphormer GROVER GraphMAE SAN GPS HGMN $(\mathcal{D})$ HGMN $(\mathcal{L})$
ogbn-arxiv 72.45 $\pm$ 1.8 71.88 $\pm$ 2.0 70.22 $\pm$ 2.3 71.50 $\pm$ 1.9 73.12 $\pm$ 1.7 74.08 $\pm$ 0.9 75.36 $\pm$ 0.7
ogbn-products 79.33 $\pm$ 1.5 78.95 $\pm$ 1.7 77.48 $\pm$ 1.8 78.02 $\pm$ 1.6 79.55 $\pm$ 1.4 80.12 $\pm$ 0.8 81.45 $\pm$ 0.6
As shown in Table 4, both variants of the proposed HGMN outperform all baseline models on the OGB datasets in terms of Micro-F1 scores. In particular, HGMN $(\mathcal{L})$ achieves the highest Micro-F1 on both ogbn-arxiv and ogbn-products, surpassing the strongest baseline GPS by 2.24% and 1.90%, respectively. Moreover, the standard deviations for the proposed models are substantially lower than those of the baselines, indicating more stable and consistent performance across multiple runs. The improvements can be attributed to HGMN’s two-stage hypergraph-based feature extraction, which effectively captures both global and local structural information, enhancing node representation learning in large-scale graphs.
Table 5: Quantitative evaluation of explainers on HGMN. HGExplainer achieves the best balance of high fidelity and low sparsity.
Explainer Dataset Fidelity (%) Sparsity (%) Stability ($\sigma$)
HGExplainer ACM 92.3 18.5 0.04
HGExplainer DBLP 89.7 21.2 0.05
GraphLIME ACM 85.1 25.4 0.07
GraphLIME DBLP 82.6 28.9 0.08
GNNExplainer ACM 87.4 22.1 0.06
GNNExplainer DBLP 84.2 24.7 0.07
Random Masking ACM 68.5 50.0 0.12
Random Masking DBLP 65.3 50.0 0.14
5.3 Interpretability Analysis
Interpretability is crucial for understanding the decision-making processes of complex models, which integrate selective state space models (e.g., Mamba) with heterogeneous graph structures. While HGMN achieves superior performance in tasks such as node classification on heterogeneous graphs, its black-box nature, stemming from metapath-based aggregations and state space selections, can obscure how predictions are formed. To address this, we conduct an interpretability analysis using post-hoc explanation methods, focusing on local explanations for individual node predictions. This reveals key metapaths and features driving HGMN’s outputs, enhancing trust and providing insights into its advantages over traditional heterogeneous GNNs.
We adopt HGExplainer, a specialized explainer for heterogeneous graph neural networks that optimizes masks over nodes, edges, and metapaths to identify important substructures while preserving semantic relations [66]. HGExplainer is post-hoc and model-agnostic, making it suitable for HGMN. It uses perturbation-based optimization: for a target node $v$ with prediction $y$, we learn masks $M_{n}$ (for node features) and $M_{e}$ (for edges/metapaths) by minimizing a loss that balances fidelity (how well the masked graph reproduces $y$) and sparsity (encouraging concise explanations). For comparison, we also evaluate GraphLIME [67], a local interpretable model explanation method for GNNs that employs HSIC Lasso to select important neighborhood features. However, GraphLIME is designed for homogeneous graphs and requires adaptations (e.g., flattening node types via one-hot encoding) for heterogeneous settings, potentially overlooking metapath semantics. We extend it by sampling metapaths as “features” but note its limitations as discussed in recent surveys on trustworthy GNNs.
We apply the explainers to HGMN trained on two benchmark heterogeneous graph datasets: ACM (3,025 papers, 5,645 authors, 1,960 subjects; node classification for papers) and DBLP (4,057 authors, 14,328 papers, 7,723 terms, 20 conferences; node classification for authors). We select 10 representative nodes per dataset from the validation set (e.g., papers in ACM classified as “database” or “machine learning”). Explanations are generated for HGMN’s predictions, with baselines including a random masking ablation and an adapted GNNExplainer for edge masking. Metrics include:
  • 1.
    Fidelity: Percentage of predictions matching the original after masking (higher is better).
  • 2.
    Sparsity: Fraction of retained elements (nodes/edges; lower is better, indicating concise explanations).
  • 3.
    Stability: Standard deviation of fidelity across 5 runs (lower is better).
HGExplainer highlights HGMN’s reliance on semantically rich metapaths. For instance, in DBLP, explaining an author’s classification as “AI researcher” reveals high importance (mask score >0.8>0.8) for author-paper-conference metapaths, where Mamba’s selective states amplify conference nodes (e.g., NeurIPS/ICLR) over less relevant terms. This underscores HGMN’s ability to route information selectively, unlike uniform aggregation in baselines. In ACM, paper classifications emphasize paper-author-subject paths, with state space selections focusing on co-author influences. GraphLIME, while effective, assigns lower fidelity in heterogeneous contexts, often overemphasizing raw features without metapath context (e.g., ignoring conference semantics in DBLP). Counterfactual analysis (e.g., masking top metapaths) shows HGMN’s predictions drop by 45% on average, confirming the explanations’ faithfulness, consistent with post-hoc methods in multi-modal heterogeneous GNNs. Table 5 summarizes the metrics averaged over 10 nodes per dataset. HGExplainer outperforms GraphLIME and baselines in fidelity and sparsity, demonstrating its suitability for HGMN.
5.4 Scalability and Efficiency Analysis
A potential challenge in integrating hypergraph convolution with state-space attention lies in the risk of increased computational overhead. To evaluate the practicality of HGMN in real-world scenarios, we conducted scalability experiments focusing on training time per epoch and peak GPU memory consumption. All models were trained on an NVIDIA A4500 GPU with 40GB memory using identical experimental settings.
Table 6 shows that, contrary to expectations, HGMN achieves lower runtime and memory usage compared to existing transformer-based graph models. In particular, both HGMN-$\mathcal{D}$ and HGMN-$\mathcal{L}$ consistently require less GPU memory and training time per epoch, while still delivering higher accuracy (Table 4). This efficiency stems from two design choices: (i) hypergraph-based aggregation reduces redundant neighborhood expansions, and (ii) the Mamba-inspired state-space attention achieves linear-time complexity, avoiding the quadratic cost of conventional attention mechanisms.
Table 6: Runtime (min/epoch) and peak GPU memory (GB) comparison on OGB datasets. Lower is better.
Model ogbn-arxiv (Mem) ogbn-products (Mem) ogbn-arxiv (Time) ogbn-products (Time)
Graphormer 28.6 GB 29.4 GB 12.5 min 18.7 min
GROVER 27.9 GB 28.7 GB 11.8 min 17.9 min
GraphMAE 26.4 GB 27.8 GB 11.2 min 16.5 min
SAN 29.1 GB 29.7 GB 13.3 min 19.4 min
GPS 28.8 GB 29.5 GB 12.9 min 18.9 min
HGMN (D) 23.2 GB 24.8 GB 8.6 min 12.1 min
HGMN (L) 22.5 GB 23.9 GB 7.9 min 11.3 min
Overall, the analysis demonstrates that HGMN provides significant performance gains (Table 4) without incurring prohibitive computational costs. The scalability of the model makes it practical for large-scale graph benchmarks, and the trade-off between efficiency and accuracy is favorable for real-world deployment.
Table 7: Ablation study of HGMN variants showing the impact of removing the residual network or Mamba block on node classification performance across multiple datasets.